Fatima's daughter is 24 years younger than Fatima.
The mean of Fatima's and Fatima's daughters age is 30.
What is Fatima's daughter's age?

Answer by stanbon(75887)

Fatima's daughter is 24 years younger than Fatima.
The mean of Fatima's and Fatima's daughters age is 30.
What is Fatima's daughter's age?
-----
Equatios:
d = F-24
(d + F-24)/2 = 30
-----
Substitute for F and solve for "d":
(d + d+24))/2 = 30
(2d+24)/2 = 30
d + 12 = 30
daughter = 18 yrs old
========================
Cheers,
Stan H.
===========